Please solve the provided system of linear equations by the elimination method.
3x=27+2y
x-7/2y=-25

Answer by TimothyLamb(4379)

3x = 27 + 2y
x + (-7/2)y = -25
---
3x + (-2)y = 27
x + (-7/2)y = -25
---
3x + (-2)y = 27
-3x + (21/2)y = 75
---
add:
---
(-2)y + (21/2)y = 27 + 75
(-4/2)y + (21/2)y = 102
(17/2)y = 102
y = 102(2/17)
---
Answer:
y = 12
x = (7/2)y - 25 = 17
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php